# Pseudo-real quantum fields

### v1

@@ Abstract
We introduce the concept of pseudo-reality for complex numbers. We show that this concept, applied to quantum fields, provides a unifying framework for two distinct approaches to pseudo-Hermitian quantum field theories. The first approach stems from analytically continuing Hermitian theories into the complex plane, while the second is based on constructing them from first principles. The pseudo-reality condition for bosonic fields resolves a long-standing problem with the formulation of gauge theories involving pseudo-Hermitian currents, sheds new light on the resolution of the so-called Hermiticity Puzzle, and leads to a consistent minimal coupling of pseudo-Hermitian quantum field theories to gravity. These results contribute significantly to the ongoing development of the first-principles construction of pseudo-Hermitian quantum field theories, including their formulation in curved spacetimes.
1 Introduction
@@
The success of non-Hermitian quantum mechanics has led to increasing interest in extending these ideas to non-Hermitian quantum field theory, with much of the progress so far focused on $PT$-symmetric quantum field theories, where the Hamiltonian is symmetric under the combined action of parity $P$ and time-reversal $T$. When the $PT$ symmetry is exact, the Hamiltonian has a real spectrum and the theory follows a (pseudo-)unitary evolution, which ensures its physical viability [1, 2, 3, 4, 5, 6] . Well-studied examples include scalar field theories with an imaginary cubic $i\phi^{3}$ [7, 8, 9, 10] or “wrong-sign” $-\phi^{4}$ quartic self-interaction [11, 12, 13].
@@
However, non-Hermitian quantum field theories bring new challenges that are not present in non-Hermitian quantum mechanics. One such issue is the Hermiticity Puzzle [14], which arises when a Hermitian theory with real fields is made non-Hermitian by introducing complex couplings. This modification leads to equations of motion being inconsistent under complex conjugation. Previous resolutions of the Hermiticity Puzzle required real fields to transform non-trivially under complex conjugation [14].
@@
Inconsistent Euler-Lagrange equations are a common feature of non-Hermitian quantum field theories, as first noted in Ref. [15], and are not limited to theories with real fields. These theories also lack Poincaré invariance, and it remains unclear how to construct consistent gauge theories or couple them to gravity. In our previous work [16] (see also Refs. [17, 18, 19]), we traced these problems to a fundamental mismatch: the field $\phi$ and its Hermitian conjugate $\phi^{\dagger}$ transform under different representations of the proper Poincaré group. This mismatch not only breaks Poincaré invariance but also causes $\phi$ and $\phi^{\dagger}$ to evolve under different Hamiltonians, $H$ and $H^{\dagger}$, respectively. To address this, we proposed replacing the Hermitian conjugate field $\phi^{\dagger}$ with a dual field $\tilde{\phi}^{\dagger}$, which transforms under the same representation of the proper Poincaré group as $\phi$ and thus, evolves with the same Hamiltonian.
@@
In this paper, we extend this idea to resolve the Hermiticity Puzzle by introducing the concept of pseudo-reality. Fields that are real in Hermitian theories become pseudo-real in pseudo-Hermitian theories, i.e., rather than being equal to their Hermitian conjugates $\phi=\phi^{\dagger}\in\mathbb{R}$, they are equal to their duals $\phi=\tilde{\phi}^{\dagger}\in\mathbb{C}$. Thus, the fields maintain the same number of degrees of freedom as a real scalar field, while being complex. This approach simultaneously resolves long-standing questions about how to gauge non-Hermitian quantum field theories and couple them to gravity in a way that is consistent with the local symmetries of the Lagrangian and current conservation. Additionally, it unifies two approaches to constructing non-Hermitian quantum field theories found in the literature: one based on the analytic continuation of Hermitian theories into the complex space, and the other based on building the non-Hermitian theory from first principles.
@@
The paper is organised as follows. In Section 2, we introduce the concept of pseudo-reality for complex functions and establish the conditions that a function must satisfy to be pseudo-real, providing examples for illustration. In Section 3, we apply the pseudo-real formalism to the non-Hermitian $i\phi^{3}$ quantum field theory. We demonstrate that not only must the field be complex, but the parity and time-reversal transformations must also be non-trivial. We find the transformation under which the Hamiltonian is pseudo-Hermitian. In Section 4, we extend this approach to higher-spin fields. We show how to consistently gauge a two complex scalar model with non-Hermitian mass mixing and outline how to couple this model to gravity.
2 Definition of pseudo-reality for complex numbers
@@
$z={\bar{z}}\qquad\ {\text{[for a real $z$]}}\,.$ (2.1)
The standard definition of reality (2.1) for a complex number $z$ involves only the operation of complex conjugation $z\to{\bar{z}}$. The notion of reality can be generalized if one considers the reality condition with respect to a complex function $f=f(z)$. To this end, we consider a complex map $f:\,\mathbb{C}\mapsto\mathbb{C}$. If the function $f=f(z)$ is analytic and satisfies the property
$f^{(-1)}(z)={\bar{f}}(z)\,,$ (2.2)
@@
$z=f({\bar{z}})\,.$ (2.3)
We then call the number $z$ a pseudo-real number with respect to the complex map $f$ ($z\in\mathbb{R}_{f}$), provided the pseudo-reality condition (2.3) is fulfilled and the function $f$ satisfies the consistency condition (2.2). Notice that Eq. (2.3) implies automatically that
$\bar{z}=f^{(-1)}(z)\;,$ (2.4)
which can also be used as a definition of the pseudo-reality, fully consistent with Eq. (2.3).
Condition (2.2) involves the inverse function $f^{(-1)}$ defined via $f^{(-1)}\Bigl{(}f(z)\Bigr{)}=z$. Therefore, the complex conjugation on the right-hand side of Eq. (2.2) acts only on the function $f$, excluding its argument $z$, i.e.,
${\bar{f}}(z)\equiv\overline{f({\bar{z}})}\,.$ (2.5)
Thus, Eq. (2.2) can be understood in terms of the Laurent series of the function $f$ and its inverse $f^{(-1)}$ as follows:
${\rm if}\quad f(z)=\sum_{n\in{\mathbb{Z}}}C_{n}z^{n}\quad{\rm and}\quad f^{(-1)}(z)=\sum_{n\in{\mathbb{Z}}}C^{(-1)}_{n}z^{n}\quad{\rm then}\quad{\overline{C}}_{n}=C^{(-1)}_{n},\ \ \forall n\in{\mathbb{Z}}\,.$ (2.6)
If one writes $z=x+iy$, then Eq. (2.3) gives us the following constraint on the variables $x$ and $y$:
$x=u(x,y)\,,\qquad y=v(x,y)\,,$ (2.7)
where $f(\bar{z})=u(x,y)+iv(x,y)$. The system of two equations (2.7) provides us, in general, with a solution in terms of a list of isolated points. In order to obtain a line, this set of generally nonlinear equations should be degenerate so that these two equations should be satisfied by a single function $y=y(x)$. It is the constraint (2.2) that makes system (2.7) degenerate.
@@
  • Take first a linear function $f(z)=C_{0}+C_{1}z$, where $C_{0}$ and $C_{1}$ are complex coefficients. In order for this function to be appropriate for the definition of the pseudo-reality of the complex number $z$ (2.3), the function $f=f(z)$ must satisfy the consistency condition (2.2). The latter condition gives us the solution for the function $f$ in the following form:
    $f(z)=e^{2i\theta}z\,,\qquad\theta\in{\mathbb{R}}\,,$ (2.8)
    thus implying that $C_{0}=0$ and $C_{1}=e^{2i\theta}$, where $\theta\in[0,2\pi)$ is a real number. Then the pseudo-reality condition (2.3) reads as follows:
    $z=e^{2i\theta}{\bar{z}}\,.$ (2.9)
    The standard reality condition (2.1) corresponds to a pseudo-real number that satisfies (2.9) with $\theta=0$. For a general angle variable $\theta$, the solution of the pseudo-reality condition (2.9) defines a straight line in the complex plane, $z=\rho e^{i\theta}$ with $\rho\in\mathbb{R}$. Alternatively, in terms of $z=x+iy$, it gives the line $y=x\tan\theta$ with the inclination of the angle $\theta$ to the abscissa. While trivial, this gives an intuitive example of what is meant by a pseudo-real number: it is one with only one degree of freedom but a non-trivial complex conjugate.
  • For a non-linear function, let us consider a generic perturbative solution in terms of $f(z)=C_{1}z+C_{2}z^{2}$, with $|C_{2}z|\ll|C_{1}|$. Solving Eq. (2.2) for small arguments, one can demonstrate that the solution is:
    $f(z)=e^{2i\theta}z\pm ige^{3i\theta}z^{2}+O(g^{2}z^{3})\,,\qquad g\in{\mathbb{R}}\,,\quad|gz|\ll 1\,.$ (2.10)
@@
    $f(z)=\frac{az+b}{cz+d}\,,\qquad{\rm with\ the\ inverse}\qquad f^{-1}(z)=\frac{dz-b}{-cz+a}\,.$ (2.11)
    The condition (2.2), which defines the allowed function $f$, gives us the following set of degenerate constraints:
    ${\overline{\mathsf{M}}}=({\rm det}\,{\mathsf{M}}){\mathsf{M}}^{-1}\,,\qquad{\mathsf{M}}=\begin{pmatrix}a&b\\ c&d\end{pmatrix}\,,$ (2.12)
    or $a=\bar{d}$, $b=-\bar{b}$, and $c=-\bar{c}$. The constraint is resolved in the form:
    ${\mathsf{M}}=\begin{pmatrix}a&i\kappa_{1}\\ i\kappa_{2}&a\end{pmatrix}\,,\qquad\kappa_{1,2}\in{\mathbb{R}}\,,\quad a\in{\mathbb{C}}\,.$ (2.13)
    Thus, the Möbius function (2.11) with the coefficients (2.13) satisfies the requirement (2.2) imposed on the function $f=f(z)$ to be used in the definition (2.2) of the pseudo-reality for a complex number $z$. Then, the linear manifold identified by Eq. (2.3) has the following explicit form:
    $(x-x_{0})^{2}+(y-y_{0})^{2}=r_{0}^{2}\,.$ (2.14)
    This formula describes a circle centered at the position $(x,y)=(x_{0},y_{0})$ with the radius $r_{0}$, related to the components of the matrix ${\mathsf{M}}$, Eq. (2.13), by the following equations, respectively:
    $x_{0}=\frac{a_{2}}{\kappa_{2}}\,,\qquad y_{0}=-\frac{a_{1}}{\kappa_{2}}\,,\qquad r_{0}^{2}=x_{0}^{2}+y^{2}_{0}-\frac{\kappa_{1}}{\kappa_{2}}\,.$ (2.15)
    Notice that, at the parameter subspace with $r^{2}_{0}<0$, no solution exists.
3 Pseudo-real $i\phi^{3}$ scalar field theory
The $i\phi^{3}$ theory, extensively studied by Bender and collaborators [7, 8, 9, 10], is the quantum field theory counterpart of the quantum mechanical model $H=p^{2}+ix^{3}$, where the spectrum was shown to be entirely real [1, 20]. Unlike the Hermitian $g\phi^{3}$ theory, which is unphysical due to its unbounded energy, the non-Hermitian $ig\phi^{3}$ theory has a bounded energy spectrum that remains real under exact $PT$ symmetry.
In this section, we revisit the $i\phi^{3}$ theory using the first-principles approach introduced in our previous work [16]. We show that extending the coupling $g\mapsto ig$ into the complex plane requires the field $\phi$ to be complex and assuming ‘standard’ parity and time-reversal transformations leads to a contradiction. To resolve this, we formulate the theory in terms of the field $\phi$ and its dual $\tilde{\phi}^{\dagger}$, both evolving with the same Hamiltonian. We introduce the concept of pseudo-reality, where $\phi$ is no longer real but a complex pseudo-real field. We look for a ‘parity-like’ transformation $\eta_{P}$ and a ‘time-reversal-like’ transformation $\eta_{T}$ with respect to which the Hamiltonian is pseudo-Hermitian and $PT$-symmetric, making this a self-consistent formulation.
@@ 3 Pseudo-real $i\phi^{3}$ scalar field theory
$\mathcal{L}(x)=\frac{1}{2}\partial_{\mu}\phi(x)\partial^{\mu}\phi(x)-\frac{1}{2}m^{2}\phi^{2}(x)-\frac{ig}{3!}\phi^{3}(x)\;,$ (3.1)
where $m^{2}>0$ and $g\in\mathbb{R}$ is some real coupling.
In the ‘standard’ approach, $\phi$ is considered to be a real scalar field, so that non-Hermiticity arises through the complex coupling $ig$:
$\mathcal{L}^{\dagger}(x)=\frac{1}{2}\partial_{\mu}\phi(x)\partial^{\mu}\phi(x)-\frac{1}{2}m^{2}\phi^{2}(x)+\frac{ig}{3!}\phi^{3}(x)\;.$ (3.2)
The non-Hermitian Lagrangian in Eq. (3.1) is $PT$-symmetric under the ‘standard’ parity and time-reversal transformations, where $\phi$ transforms as a pseudo-scalar field:
$P:$ $\>\>\>\>\phi(x)\longmapsto\phi^{P}(x_{P})=-\phi(x)\>,\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\hat{\mathcal{P}}\hat{\phi}(x_{P})\hat{\mathcal{P}}^{-1}=-\hat{\phi}(x)\>,$ (3.3a)
$T:$ $\>\>\>\>\phi(x)\longmapsto\phi^{T}(x_{T})=\phi(x)\>,\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\hat{\mathcal{T}}\hat{\phi}(x_{T})\hat{\mathcal{T}}^{-1}=\hat{\phi}(x)\>,$ (3.3b)
$PT:$ $\>\>\mathscr{L}(x)\longmapsto\mathscr{L}^{PT}(x_{PT})=\mathscr{L}(x)\>,\>\>\>\>\>\>\>\>\>\>(\hat{\mathcal{P}}\hat{\mathcal{T}})\hat{\mathscr{L}}(x_{PT})(\hat{\mathcal{P}}\hat{\mathcal{T}})^{-1}=\hat{\mathscr{L}}(x)\>.$ (3.3c)
Here, we consider the transformation of the field function $\phi$ on the left and the field operator $\hat{\phi}$ on the right.
The existence of an anti-linear symmetry implies that the Lagrangian is pseudo-Hermitian [6, 21], i.e., there exists some linear transformation $\eta$ such that it takes the Lagrangian to its Hermitian conjugate:
$\eta:\>\mathscr{L}(x)\longmapsto\mathscr{L}^{\eta}(x_{\eta})=\mathscr{L}^{\dagger}(x)\>,\>\>\>\>\>\>\hat{\eta}\hat{\mathscr{L}}(x_{\eta})\hat{\eta}^{-1}=\hat{\mathscr{L}}^{\dagger}(x)\>.$ (3.4)
Here, we take into account that $\eta:x\mapsto x_{\eta}$ could be a coordinate transformation (e.g., parity, $\mathcal{P}:x\mapsto x_{P}$).
Indeed, the Lagrangian in Eq. (3.1) is pseudo-Hermitian with respect to the ‘standard’ parity transformation in Eq. (3.3a):
$P:\>\mathscr{L}(x)\longmapsto\mathscr{L}^{P}(x_{P})=\mathscr{L}^{\dagger}(x)\>,\text{\>\>\>\>\>\>\>}\hat{\mathcal{P}}\hat{\mathscr{L}}(x_{P})\hat{\mathcal{P}}^{-1}=\hat{\mathscr{L}}^{\dagger}(x)\>.$ (3.5)
It is also anti-pseudo-Hermitian with respect to the ‘standard’ time-reversal transformation in Eq. (3.3b):
$T:\>\mathscr{L}(x)\longmapsto\mathscr{L}^{T}(x_{T})=\mathscr{L}^{\dagger}(x)\>,\>\>\>\>\>\>\>\>\>\hat{\mathcal{T}}\hat{\mathscr{L}}(x_{T})\hat{\mathcal{T}}^{-1}=\hat{\mathscr{L}}^{\dagger}(x)\>.$ (3.6)
We refer to the system satisfying Eqs. (3.5) and (3.6) as possessing the property of ‘anti-pseudo-Hermiticity’, since the time-reversal operator $T$ is anti-linear [6]. The corresponding Hamiltonian
$H=\int{\rm d}^{3}\vec{x}\>\quantity(\frac{1}{2}\dot{\phi}^{2}+\frac{1}{2}(\nabla\phi)^{2}+\frac{1}{2}m^{2}\phi^{2}+\frac{ig}{3!}\phi^{3})$ (3.7)
@@
$P:\>$ $H\longmapsto H^{\dagger}\>,\text{\>\>\>\>\>\>}\>\>\>\hat{\mathcal{P}}\hat{H}\hat{\mathcal{P}}^{-1}=\hat{H}^{\dagger}\>,$ (3.8a)
$T:\>$ $H\longmapsto H^{\dagger}\>,\text{\>\>\>\>\>\>\>}\>\>\hat{\mathcal{T}}\hat{H}\hat{\mathcal{T}}^{-1}=\hat{H}^{\dagger}\>.$ (3.8b)
3.1 Issues with the standard approach

### v2

@@ Abstract
We introduce the concept of pseudo-reality for complex numbers. We show that this concept, applied to quantum fields, provides a unifying framework for two distinct approaches to pseudo-Hermitian quantum field theories. The first approach stems from analytically continuing Hermitian theories into the complex plane, while the second is based on constructing them from first principles. The pseudo-reality condition for bosonic fields resolves a long-standing problem with the formulation of gauge theories involving pseudo-Hermitian currents, sheds new light on the resolution of the so-called Hermiticity Puzzle, and may allow a consistent minimal coupling of pseudo-Hermitian quantum field theories to gravity. We focus on the $i\phi^{3}$ cubic scalar theory, obtaining the relevant pseudo-reality conditions up to quadratic order in the coupling; a theory of two complex scalar fields with non-Hermitian mass mixing; and the latter’s coupling to a $U(1)$ gauge field. The general principle of pseudo-reality, however, is expected to contribute to the ongoing development of the first-principles construction of pseudo-Hermitian quantum field theories, including their formulation in curved spacetimes.
This is an author-prepared post-print of Phys. Rev. D 112 (2025) 065007, published by the American Physical Society under the terms of the CC BY 4.0 license (funded by SCOAP3).
I Introduction
@@
The success of non-Hermitian quantum mechanics has led to increasing interest in extending these ideas to non-Hermitian quantum field theory, with much of the progress so far focused on $PT$-symmetric quantum field theories, where the Hamiltonian is symmetric under the combined action of parity $P$ and time-reversal $T$. When the $PT$ symmetry is exact, the Hamiltonian has a real spectrum and the theory follows a (pseudo-)unitary evolution, which ensures its physical viability Bender:1 ; Bender:2 ; Bender:2009mq ; Most:1 ; Most:2 ; Most:3 . Well-studied examples include scalar field theories with an imaginary cubic $i\phi^{3}$ Bphi1 ; Bphi2 ; Bphi3 ; Bphi4 or “wrong-sign” $\phi^{4}$ quartic self-interaction Sphi1 ; Sphi2 ; Bphi5 .
@@
However, non-Hermitian quantum field theories bring new challenges that are not present in non-Hermitian quantum mechanics. One such issue is the Hermiticity Puzzle Mannheim:2018dur , which arises when a Hermitian theory with real fields is made non-Hermitian by introducing complex couplings. This modification leads to equations of motion that are inconsistent under complex conjugation. Previous resolutions of the Hermiticity Puzzle required real fields to transform non-trivially under complex conjugation Mannheim:2018dur .
@@
Inconsistent Euler-Lagrange equations are a common feature of non-Hermitian quantum field theories, as first noted in Ref. Alexandre:2017foi , and are not limited to theories with real fields. These theories also lack Poincaré invariance, and it remains unclear how to construct consistent gauge theories or couple them to gravity. In our previous work Sablevice:2023odu (see also Refs. Alexandre:2020gah ; Alexandre:2022uns ; Chernodub:2021waz ), we traced these problems to a fundamental mismatch: the field $\phi$ and its Hermitian conjugate $\phi^{\dagger}$ transform under different representations of the proper Poincaré group. This mismatch not only breaks Poincaré invariance but also causes $\phi$ and $\phi^{\dagger}$ to evolve under different Hamiltonians, $H$ and $H^{\dagger}$, respectively. To address this, we proposed replacing the Hermitian conjugate field $\phi^{\dagger}$ with a dual field $\tilde{\phi}^{\dagger}$, which transforms under the same representation of the proper Poincaré group as $\phi$ and thus evolves with the same Hamiltonian.
@@
In this paper, we extend this idea to resolve the Hermiticity Puzzle by introducing the concept of pseudo-reality. Fields that are real in Hermitian theories become pseudo-real in pseudo-Hermitian theories, i.e., rather than being equal to their Hermitian conjugates $\phi=\phi^{\dagger}\in\mathbb{R}$, they are equal to their duals $\phi=\tilde{\phi}^{\dagger}\in\mathbb{C}$. Thus, the fields maintain the same number of degrees of freedom as a real scalar field, while being complex. This approach resolves a long-standing question about how to gauge non-Hermitian quantum field theories and may allow to couple them to gravity in a way that is consistent with the local symmetries of the Lagrangian and current conservation. Additionally, it unifies two approaches to constructing non-Hermitian quantum field theories found in the literature: one based on the analytic continuation of Hermitian theories into the complex space, and the other based on building the non-Hermitian theory from first principles.
@@
The paper is organised as follows. In Section 2, we introduce the concept of pseudo-reality for complex functions and establish the conditions that a function must satisfy to be pseudo-real, providing examples for illustration. In Section 3, we apply the pseudo-real formalism to the non-Hermitian $i\phi^{3}$ quantum field theory. We demonstrate that not only must the field be complex, but the parity and time-reversal transformations must also be non-trivial. We find the transformation under which the Hamiltonian is pseudo-Hermitian. In Section 4, we extend this approach to higher-spin fields. We show how to consistently gauge a two complex scalar model with non-Hermitian mass mixing and conjecture how to couple this model to gravity.
II Definition of pseudo-reality for complex numbers
@@
[…]
The standard definition of reality (1) for a complex number $z$ involves only the operation of complex conjugation $z\to{\bar{z}}$. The notion of reality can be generalized if one considers the reality condition with respect to a complex function $f=f(z)$. To this end, we consider a complex map $f:\,\mathbb{C}\mapsto\mathbb{C}$. If the function $f=f(z)$ is analytic and satisfies the property
$f^{(-1)}(z)={\bar{f}}(z)\,$, (2)
@@
$z=f({\bar{z}})\,.$ (3)
We then call the number $z$ a pseudo-real number with respect to the complex map $f$ ($z\in\mathbb{R}_{f}$), provided the pseudo-reality condition (3) is fulfilled and the function $f$ satisfies the consistency condition (2). Notice that Eq. (3) implies automatically that
$\bar{z}=f^{(-1)}(z)\;,$ (4)
which can also be used as a definition of the pseudo-reality, fully consistent with Eq. (3).
Condition (2) involves the inverse function $f^{(-1)}$ defined via $f^{(-1)}\Bigl(f(z)\Bigr)=z$. Therefore, the complex conjugation on the right-hand side of Eq. (2) acts only on the function $f$, excluding its argument $z$, i.e.,
${\bar{f}}(z)\equiv\overline{f({\bar{z}})}\,.$ (5)
Thus, Eq. (2) can be understood in terms of the Laurent series of the function $f$ and its inverse $f^{(-1)}$ as follows:
${\rm if}\quad f(z)=\sum_{n\in{\mathbb{Z}}}C_{n}z^{n}\quad{\rm and}\quad f^{(-1)}(z)=\sum_{n\in{\mathbb{Z}}}C^{(-1)}_{n}z^{n}\quad{\rm then}\quad{\overline{C}}_{n}=C^{(-1)}_{n},\ \ \forall n\in{\mathbb{Z}}\,.$ (6)
If one writes $z=x+iy$, then Eq. (3) gives us the following constraints on the variables $x$ and $y$:
$x=u(x,y)\,,\qquad y=v(x,y)\,$, (7)
where $f(\bar{z})=u(x,y)+iv(x,y)$. The system of two equations (7) provides us, in general, with a solution in terms of a list of isolated points. In order to obtain a line, this set of generally nonlinear equations should be degenerate so that these two equations should be satisfied by a single function $y=y(x)$. It is the constraint (2) that makes the system (7) degenerate.
@@
  • Take first a linear function $f(z)=C_{0}+C_{1}z$, where $C_{0}$ and $C_{1}$ are complex coefficients. In order for this function to be appropriate for the definition of the pseudo-reality of the complex number $z$ (3), the function $f=f(z)$ must satisfy the consistency condition (2). The latter condition gives us the solution for the function $f$ in the following form:
    $f(z)=e^{2i\theta}z\,,\qquad\theta\in{\mathbb{R}}\,$, (8)
    thus implying that $C_{0}=0$ and $C_{1}=e^{2i\theta}$, where $\theta\in[0,2\pi)$ is a real number. Then the pseudo-reality condition (3) reads as follows:
    $z=e^{2i\theta}{\bar{z}}\,.$ (9)
    The standard reality condition (1) corresponds to a pseudo-real number that satisfies (9) with $\theta=0$. For a general angle variable $\theta$, the solution of the pseudo-reality condition (9) defines a straight line in the complex plane, $z=\rho e^{i\theta}$ with $\rho\in\mathbb{R}$. Alternatively, in terms of $z=x+iy$, it gives the line $y=x\tan\theta$ with the inclination of the angle $\theta$ to the abscissa. While trivial, this gives an intuitive example of what is meant by a pseudo-real number: it is one with only one degree of freedom but a non-trivial complex conjugate.
  • For a non-linear function, let us consider a generic perturbative solution in terms of $f(z)=C_{1}z+C_{2}z^{2}$, with $|C_{2}z|\ll|C_{1}|$. Solving Eq. (2) for small arguments, one can demonstrate that the solution is:
    $f(z)=e^{2i\theta}z\pm ige^{3i\theta}z^{2}+O(g^{2}z^{3})\,,\qquad g\in{\mathbb{R}}\,,\quad|gz|\ll 1\,.$ (10)
@@
    $f(z)=\frac{az+b}{cz+d}\,$, (11)
    with the inverse
    $f^{-1}(z)=\frac{dz-b}{-cz+a}\,.$ (12)
    The condition (2), which defines the allowed function $f$, gives us the following set of degenerate constraints:
    ${\overline{\mathsf{M}}}=({\rm det}\,{\mathsf{M}}){\mathsf{M}}^{-1}\,,\qquad{\mathsf{M}}=\begin{pmatrix}a&b\\ c&d\end{pmatrix}\,$, (13)
    or $a=\bar{d}$, $b=-\bar{b}$, and $c=-\bar{c}$. The constraint is resolved in the form
    ${\mathsf{M}}=\begin{pmatrix}a&i\kappa_{1}\\ i\kappa_{2}&a\end{pmatrix}\,,\qquad\kappa_{1,2}\in{\mathbb{R}}\,,\quad a\in{\mathbb{C}}\,.$ (14)
    Thus, the Möbius function (11) with the coefficients (14) satisfies the requirement (2) imposed on the function $f=f(z)$ to be used in the definition (2) of the pseudo-reality for a complex number $z$. Then, the linear manifold identified by Eq. (3) has the following explicit form:
    $(x-x_{0})^{2}+(y-y_{0})^{2}=r_{0}^{2}\,.$ (15)
    This formula describes a circle centered at the position $(x,y)=(x_{0},y_{0})$ with the radius $r_{0}$, related to the components of the matrix $\mathsf{M}$}, Eq. (14), by the following equations:
    $x_{0}=\frac{a_{2}}{\kappa_{2}}\,,\qquad y_{0}=-\frac{a_{1}}{\kappa_{2}}\,,\qquad r_{0}^{2}=x_{0}^{2}+y^{2}_{0}-\frac{\kappa_{1}}{\kappa_{2}}\,.$ (16)
    Notice that, in the region of parameter space with $r^{2}_{0}<0$, no solution exists.
In the sections that follow, we will apply this concept of pseudo-reality to quantum fields.
III Pseudo-real $i\phi^{3}$ scalar field theory
The $i\phi^{3}$ theory, extensively studied by Bender and collaborators Bphi1 ; Bphi2 ; Bphi3 ; Bphi4 , is the quantum field theory counterpart of the quantum mechanical model $H=p^{2}+ix^{3}$, for which the spectrum was shown to be entirely real Bender:1 ; Bender:2012yv . Unlike the Hermitian $g\phi^{3}$ theory, which is unphysical due to its unbounded energy, the non-Hermitian $ig\phi^{3}$ theory has a bounded energy spectrum that remains real under exact $PT$ symmetry.
In this section, we revisit the $i\phi^{3}$ theory using the first-principles approach introduced in previous work Sablevice:2023odu . We show that extending the coupling $g\mapsto ig$ into the complex plane requires the field $\phi$ to be complex and the ‘standard’ parity and time-reversal transformations lead to a contradiction. To resolve this, we formulate the theory in terms of the field $\phi$ and its dual $\tilde{\phi}^{\dagger}$, both evolving with the same Hamiltonian. We apply the concept of pseudo-reality, where $\phi$ is no longer real but a complex pseudo-real field, and we look for a ‘parity-like’ transformation $\eta_{P}$ and a ‘time-reversal-like’ transformation $\eta_{T}$ with respect to which the Hamiltonian is pseudo-Hermitian and $PT$ symmetric, making this a self-consistent formulation.
@@ III Pseudo-real $i\phi^{3}$ scalar field theory
$\mathcal{L}(x)=\frac{1}{2}\partial_{\mu}\phi(x)\partial^{\mu}\phi(x)-\frac{1}{2}m^{2}\phi^{2}(x)-\frac{ig}{3!}\phi^{3}(x)\;,$ (17)
where $m^{2}>0$ and $g\in\mathbb{R}$ is some real coupling. In the ‘standard’ approach, $\phi$ is considered to be a real scalar field, so that non-Hermiticity arises through the complex coupling $ig$, with
$\mathcal{L}^{\dagger}(x)=\frac{1}{2}\partial_{\mu}\phi(x)\partial^{\mu}\phi(x)-\frac{1}{2}m^{2}\phi^{2}(x)+\frac{ig}{3!}\phi^{3}(x)\;.$ (18)
The non-Hermitian Lagrangian in Eq. (17) is $PT$ symmetric under the ‘standard’ parity and time-reversal transformations, where $\phi$ transforms as a pseudo-scalar field:
$P:$ $\>\>\>\>\phi(x)\longmapsto\phi^{P}(x_{P})=-\phi(x)\>,\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\hat{\mathcal{P}}\hat{\phi}(x_{P})\hat{\mathcal{P}}^{-1}=-\hat{\phi}(x)\>,$ (19a)
$T:$ $\>\>\>\>\phi(x)\longmapsto\phi^{T}(x_{T})=\phi(x)\>,\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>\hat{\mathcal{T}}\hat{\phi}(x_{T})\hat{\mathcal{T}}^{-1}=\hat{\phi}(x)\>,$ (19b)
$PT:$ $\>\>\mathscr{L}(x)\longmapsto\mathscr{L}^{PT}(x_{PT})=\mathscr{L}(x)\>,\>\>\>\>\>\>\>\>\>\>(\hat{\mathcal{P}}\hat{\mathcal{T}})\hat{\mathscr{L}}(x_{PT})(\hat{\mathcal{P}}\hat{\mathcal{T}})^{-1}=\hat{\mathscr{L}}(x)\>.$ (19c)
Here, the transformation of the $c$-number field $\phi$ is given on the left, and the transformation of the field operator $\hat{\phi}$, indicated by a caret, on the right.
The existence of an anti-linear symmetry implies that the Lagrangian is pseudo-Hermitian Most:3 ; Most:4 , i.e., there exists some linear transformation $\eta$ that takes the Lagrangian to its Hermitian conjugate:
$\eta:\>\mathscr{L}(x)\longmapsto\mathscr{L}^{\eta}(x_{\eta})=\mathscr{L}^{\dagger}(x)\>,\>\>\>\>\>\>\hat{\eta}\hat{\mathscr{L}}(x_{\eta})\hat{\eta}^{-1}=\hat{\mathscr{L}}^{\dagger}(x)\>.$ (20)
Here, we take into account that $\eta:x\mapsto x_{\eta}$ could be a coordinate transformation (e.g., parity, $P:x\mapsto x_{P}$).
Indeed, the Lagrangian in Eq. (17) is pseudo-Hermitian with respect to the ‘standard’ parity transformation in Eq. (19a):
$P:\>\mathscr{L}(x)\longmapsto\mathscr{L}^{P}(x_{P})=\mathscr{L}^{\dagger}(x)\>,\text{\>\>\>\>\>\>\>}\hat{\mathcal{P}}\hat{\mathscr{L}}(x_{P})\hat{\mathcal{P}}^{-1}=\hat{\mathscr{L}}^{\dagger}(x)\>.$ (21)
It is also anti-pseudo-Hermitian with respect to the ‘standard’ time-reversal transformation in Eq. (19b):
$T:\>\mathscr{L}(x)\longmapsto\mathscr{L}^{T}(x_{T})=\mathscr{L}^{\dagger}(x)\>,\>\>\>\>\>\>\>\>\>\hat{\mathcal{T}}\hat{\mathscr{L}}(x_{T})\hat{\mathcal{T}}^{-1}=\hat{\mathscr{L}}^{\dagger}(x)\>.$ (22)
We call this anti-pseudo-Hermiticity due to the time-reversal operator being anti-linear Most:3 . The corresponding Hamiltonian
$H=\int{\rm d}^{3}\vec{x}\>\quantity(\frac{1}{2}\dot{\phi}^{2}(t,\vec{x})+\frac{1}{2}(\nabla\phi(t,\vec{x}))^{2}+\frac{1}{2}m^{2}\phi^{2}(t,\vec{x})+\frac{ig}{3!}\phi^{3}(t,\vec{x}))$ (23)
@@
$P:\>$ $H\longmapsto H^{\dagger}\>,\text{\>\>\>\>\>\>}\>\>\>\hat{\mathcal{P}}\hat{H}\hat{\mathcal{P}}^{-1}=\hat{H}^{\dagger}\>,$ (24a)
$T:\>$ $H\longmapsto H^{\dagger}\>,\text{\>\>\>\>\>\>\>}\>\>\hat{\mathcal{T}}\hat{H}\hat{\mathcal{T}}^{-1}=\hat{H}^{\dagger}\>.$ (24b)
III.1 Issues with the standard approach